OpenStudy (anonymous):

write the exponential function (y=a*b^x) that passes through (0,30) and (2,750).

geerky42 (geerky42):

You can plug in pairs like this:\[30 = a\cdot b^0\]\[750 = a\cdot b^2\] Can you solve for a and b?

OpenStudy (anonymous):

a=30 and b=5?

geerky42 (geerky42):

That's right.

OpenStudy (anonymous):

that was easier than i expected haha

geerky42 (geerky42):

yeah lol.